OpenStudy (blondisaurus):

what is the equation of the line that passes through (3,4) and (-1,5)?

OpenStudy (blondisaurus):

I dont get how they arranged it to equal x+4y=19.

OpenStudy (anonymous):

find slope (y2-y1)/(x2-x1)

OpenStudy (blondisaurus):

-1/4.

OpenStudy (anonymous):

y-y2=m(x-x2) this is the equation plug in

OpenStudy (anonymous):

you could also use y1 instead of y2

OpenStudy (blondisaurus):

I got y= -1/4x-3.25

OpenStudy (anonymous):

y+3.25=-1/4 x

OpenStudy (anonymous):

4y+13=-x 4y+x=-13

OpenStudy (anonymous):

i think you got -3.25 wrong or question you posted is a bit wrong

OpenStudy (blondisaurus):

Yes, it was 4.75. not 3.25. Thank you!